Given this information for a monopoly company producing cars :


  1. cost of production : 2 million ( 2000 in thousands) w/ nothing to maintain 
  2. foreseeable demand : q= 900-150p ( q is the number of drivers (thousands) given the price per driver p)


What would be the profit-maximizing problem (optimization)


What would be the profit-maximizing quantity and price?



1
Expert's answer
2021-02-22T14:03:03-0500

Solution:

Profit maximizing problem is given by:

P(q) = R(q) – C(q)

Where: P(q) = Profit

R(q) = Total Revenue (TR)

C(q) = Total Cost (TC)


Given the demand function: q = 900 – 150p, compute the inverse demand function by solving for p in the demand function.

q = 900 – 150p

p = "(\frac{q-900 }{150} )" = 0.0067q – 6


Find TR = "P\times Q"

="(0.0067q - 6) (Q)"

= "0.0067Q^{2} -6Q" 0.0067Q2 – 6Q


Derive the Marginal Revenue (MR) by getting the derivative of the TR function:

MR = 0.0134Q – 6

Find the Total Cost (TC) function = C x Q

= "(2,000,000) (Q)" = 2,000,000Q


Derive the Marginal Cost (MC) by getting the derivative of the TC function:

MC = 2,000,000

Profit Maximization is where: MR = MC

= 0.0134Q – 6 = 2,000,000

= 0.0134Q = 2,000,000 + 6

= 0.0134Q = 2,000,006


Q ="\frac{2,000,006}{0.0134}"


Q = 149,254,179


The profit maximizing quantity is = 149,254,179 units


Find P = 0.0067Q – 6

Substitute for Q:

= "(0.0067\times 149,254,178) - 6"

= 1000,003 – 6


P = 999,997


The profit-maximizing price = 999,997